A 23 kg object is pushed up an inclined plane with angle of inclination of 14 degrees. The coefficient of kinetic friction between the object and the incline is 0.30. Determine the acceleration of the object if it is pushed by a 250 N force that is acting parallel to the incline.


1
Expert's answer
2022-05-26T14:07:28-0400
"a=\frac{F-F_f-mg\sin\theta}{m}""F_f=\mu N=\mu mg\cos\theta""a=\frac{F-\mu mg\cos\theta-mg\sin\theta}{m}\\
=F/m-\mu g \cos\theta-g\sin\theta""a=250/23-0.30*9.8 \cos14^\circ-9.8\sin14^\circ\\
=5.6\:\rm m/s^2"